# A Dataless Reinforcement Learning Approach to Rounding Hyperplane Optimization for Max-Cut

### v1

@@ 4.1 Settings
Table 2: Average and maximum cut values for the ER dataset obtained from a batch of 512 hyperplanes of pGW against the proposed Agent. All graphs here have a total of 1000 nodes and are ER graphs generated from NetworkX Package. We observe that the proposed agent is able to get a higher score than pGW on all seeds across all densities as indicated by the probability of edge creation (column 2).
@@
For baselines, we use the pGW in Algorithm 2.1 where we sample a batch of 512 unit norm vectors uniformly in parallel. We apply each of these vectors to (5) to obtain the value of Cut for a given graph. We report the average of these values as well as the maximum.
@@ Appendix A Impact of the number of parameters in our agent’s NN
Here, we include ablation studies on randomly generated ER graphs to determine if having higher number of learnable parameters in the encoder layer is beneficial to seeing higher cuts. To this end, we present results across different seeds, densities and number of hidden neurons in our NN, i.e., $l$. We ran ablation studies for 2000 time steps. From these runs we see that having a higher number of learnable parameters generally leads to higher Cut values than pGW.
@@
Table 3: Results of our agent using different number of hidden dimension, $l$ for ER graph with $p=0.29$ and seed $=1$. We see that the average and the maximum cut from agent have an increasing trend with increase in hidden dimension.
@@
Table 4: Results of our agent using different number of hidden dimension, $l$ for ER graph with $p=0.85$ and seed $=1$. We see that the average and the maximum cut from agent have an increasing trend with increase in hidden dimension.
@@
Table 5: Results of our agent using different number of hidden dimension, $l$ for ER graph with $p=0.29$ and seed $=50$. We see that the average and the maximum cut from agent have an increasing trend with increase in hidden dimension.
@@
Table 6: Results of our agent using different number of hidden dimension, $l$ for ER graph with $p=0.85$ and seed $=50$. We see that the average and the maximum cut from agent have an increasing trend with increase in hidden dimension.

### v2

@@ 4.1 Settings
Table 2: Average and maximum cut values for the ER dataset obtained from a batch of 256 hyperplanes of pGW against the proposed Agent. All graphs here have a total of 1000 nodes and are ER graphs generated from NetworkX Package. We observe that the proposed agent is able to get a higher score than pGW on all seeds across all densities as indicated by the probability of edge creation (column 2).
@@
For baselines, we use the pGW in Algorithm 2.1 where we sample a batch of 256 unit norm vectors uniformly in parallel. We apply each of these vectors to (5) to obtain the value of Cut for a given graph. We report the average of these values as well as the maximum.
@@ Appendix A Impact of the number of parameters in our agent’s NN
Here, we include ablation studies on randomly generated ER graphs to determine if having higher number of learnable parameters in the encoder layer is beneficial to seeing higher cuts. To this end, we present results across different seeds, densities and number of hidden neurons in our NN, i.e., $l$. We ran ablation studies for 2000 time steps. We see that the average cut increased with increase in $l$ until $l=512$ and then decreased in Tables 3, 4, and, 6 whereas in 5 the average cut increased with higher values of $l$.
@@
Table 3: Results of our agent using different number of hidden dimension, $l$ for ER graph with $p=0.29$ and seed $=1$. We see that the average cut generally increases until $l=512$.
@@
Table 4: Results of our agent using different number of hidden dimension, $l$ for ER graph with $p=0.85$ and seed $=1$. We see that the average cut increases until $l=512$ and then decreases.
@@
Table 5: Results of our agent using different number of hidden dimension, $l$ for ER graph with $p=0.29$ and seed $=50$. We see that the average cut from agent have an increasing trend with increase in hidden dimension.
@@
Table 6: Results of our agent using different number of hidden dimension, $l$ for ER graph with $p=0.85$ and seed $=50$. We see that the average cut from agent have an increasing trend with increase in hidden dimension until $l=512$ after which it decreases slightly.